OpenStudy (anonymous):

Help !! Will Fan and Medall!!:)) Find the GCF from the following Polynomials 18x^4+9x^3-27x+3

OpenStudy (anonymous):

\[18x^4=2*3*3*x*x*x*x\] \[9x^3=3*3*x*x*x\] \[27x=3*3*3*x\] take 3*3*x common in that equation

OpenStudy (anonymous):

24x^4-6x^3+12x^2-18x @aryandecoolest ??:))

OpenStudy (anonymous):

 do this thing for rest of the polynomial, just factor it to the lowest possible prime numbers and take out common terms from there hope u can @JayBaybie